# CogDevelop2K: Reversed Cognitive Development in Multimodal Large Language Models

### v1

@@ Abstract
Are Multi-modal Large Language Models (MLLMs) stochastic parrots? Do they genuinely understand and are capable of performing the tasks they excel at? This paper aims to explore the fundamental basis of MLLMs, i.e. core cognitive abilities that human intelligence builds upon to perceive, comprehend, and reason. To this end, we propose CogDevelop2K, a comprehensive benchmark that spans 12 sub-concepts from fundamental knowledge like object permanence and boundary to advanced reasoning like intentionality understanding, structured via the developmental trajectory of a human mind. We evaluate 46 MLLMs on our benchmarks. Comprehensively, we further evaluate the influence of evaluation strategies and prompting techniques. Surprisingly, we observe a reversed cognitive developmental trajectory compared to humans.
@@ 1 Introduction
Building on the foundation of advanced large language models (LLMs), multi-modal large language models (MLLMs) have recently demonstrated human-level performance in complex tasks involving high-level reasoning, perception, and cognition Li et al. (2024a); Liu et al. (2024); Team (2023); Fu et al. (2023); OpenAI (2023) such as Spatial Reasoning Chen et al. (2024); Cai et al. (2024), OCR Mori et al. (1999), Scene Understanding Cordts et al. (2016); Chen et al. (2017), Action Recognition Jhuang et al. (2013); Herath et al. (2017) and Prediction Lan et al. (2014); Kong & Fu (2022), etc. The progress in MLLMs has reignited hopes for achieving Artificial General Intelligence (AGI). However, we pose a critical question: Do MLLMs truly comprehend these tasks and possess the genuine capabilities to perform them, or are they merely ”stochastic parrots” that rely on learning spurious correlations? To explore this, we draw inspiration from the development of human cognition.
@@
As highlighted in previous research, core cognitive abilities form the foundational basis of higher-level human intelligence that existing MLLMs excel at, but such excellency do not translate into a general domain (Mitchell, 2020; 2021; Shiffrin & Mitchell, 2023). The performance of MLLMs on core cognitive tasks therefore provides a more profound insight into their true capacities for knowledge, reasoning, and perception. This, in turn, serves as a critical indicator of whether MLLMs possess genuine intelligence or if they are merely ”stochastic parrots” that depend on learning spurious correlations. To this end, we draw on theoretical and empirical approaches from developmental science to create benchmarks that evaluate core cognitive abilities in large vision-language models, examining the relationships between these abilities. On a high level, we follow Jean Piaget’s theory of cognitive development, which identifies four stages in children: sensorimotor, preoperational, concrete operational, and formal operational (Piaget, 1950; Piaget & Inhelder, 1969; 1974). During the sensorimotor stage, infants acquire knowledge through sensory experiences and actions, developing an understanding of basic object properties, such as permanence, continuity, and boundaries. In the preoperational stage, symbolic representation emerges, along with a grasp of basic physical properties. The concrete operational stage is characterized by the development of logical thinking and an understanding of intuitive physics. Finally, the formal operational stage introduces more advanced cognitive abilities, including abstraction, hypothetical reasoning, counterfactual thinking, and tool use. The interdependence and developmental trajectories of these abilities can be mapped in terms of a tree-like structure (as illustrated in Fig. 1).
To evaluate the performance of MLLMs on the core cognitive abilities, we curate the first-ever vision cognitive development benchmark, termed as CogDeveop2K, which consists of a total of 2519 questions with 2517 images and 455 videos. Then, we evaluate 46 MLLM models on our benchmark that spans all four stages of cognitive development. We introduce a novel multi-frame question format to evaluate models’ co-reference, cognitive reasoning and temporal understanding capability simultaneously. Forty-seven models are compared against a human baseline under zero-shot conditions using 11 different prompts (including no prompt). Surprisingly, while prompts can boost model performance by 8.1%, models still demonstrate reversed trends in cognitive development against those observed in children. For example, GPT series perform better in formal operation stage while performing worse in concrete operation stage.
@@ 3.1 Multi-modal Large Language Models
The Vision Language Model (VLM) has a long history from Convolution Neural Networks (CNN) and Recurrent Neural Networks (RNN) (Karpathy & Fei-Fei, 2014; Vinyals et al., 2015) to unified modeling of visual and text modality with transformers (Li et al., 2019; Xu et al., 2023; Tan & Bansal, 2019; Alayrac et al., 2022; Radford et al., 2021). With the advancement of Large Language Models (LLMs), existing state-of-the-art MLLMs (Liu et al., 2024; Li et al., 2023) adopt open-sourced Large Language Models such as Llama (Touvron et al., 2023), Mistral (Jiang et al., 2023), etc. Instruction Tuning is also introduced to further improve the task generalization ability of MLLMs (Liu et al., 2024; Dai et al., 2023). To acquire open-ended conversation abilities, LLaVA (Liu et al., 2024) proposes to distill the conversational abilities of ChatGPT to MLLMs, boosting performance by a large margin, which becomes a defacto procedure in the area (Wang et al., 2023; Bai et al., 2023; Team, 2023; 2024; Sun et al., 2023; Li et al., 2022).
@@ 3.2 Human Cognitive Development
Preoperational and concrete operational stage are the second and third stage of Piaget’s cognitive development. Typically spanning over 2 to 7 years of age, preoperational stage is the transitional stage to concrete operational stage, which children enters around 7 years of age. During then, children begin to develop internalized mental actions supported by organized structures that can be manipulated and reversed in systematic ways, known as mental operatios (Janet, 1905; Kirkpatrick, 1908; Piaget, 1950; Piaget & Inhelder, 2014; Miller, 2016). Through mental operations, children are then able to rigidly perform tasks that are previously unreachable, such as thinking from other people’s perspectives, understanding hierarchical relations of objects, and reasoning about physical events in the world. These tasks require not only rudimentary understandings of physical concepts, which gradually became in place during preoperational stage, but also relational and transformational reasoning that can only be done through mental operations (Piaget & Inhelder, 1974; Church & Goldin-Meadow, 1986; Houdé, 1997). Since preoperational stage is mostly meaningful as the transitional period preceding concrete operational stage, we do not have evaluation dimensions specifically targeting the stage. However, tasks targeting concrete operational stage could assess presentations of knowledge associated with preoperational stage, as prominently illustrated by the law of conservation (Piaget, 1952; Halford, 2011; Houdé, 1997).
@@ 3.3 Evaluation Dimension
Boundary refers to the cognitive recognition of where one object ends and another begins, an essential aspect of perceiving and understanding the physical world (Kestenbaum et al., 1987). Without understanding boundary, which means where the object ends, it seems very hard to construct a concept of object (Berkeley, 1709; Jackendoff, 1991).
@@
Permanence, or specifically object permanence, is the idea in cognitive development where an individual understands that objects continue to exist even when they are not visible (Baillargeon, 1986; Spelke et al., 1992). Imagine a simple scene: a small child playing peek-a-boo. In the beginning, when the caregiver covers their face with their hands, the child might seem surprised or even distressed, thinking the person has disappeared. However, as the child’s understanding of permanence develops, they begin to realize that just because they can’t see the person’s face, it doesn’t mean the person is gone.
@@
Hierarchical relation refers to the cognitive phenomena that children begin to understand hierarchical relations and be able to organize objects or concepts into structured categories and subcategories, which are supported by the development of mental operations marked by class inclusion and transitivity (Shipley, 1979; Winer, 1980; Chapman & McBride, 1992). Class inclusion refers to the ability to recognize that some classes or groups of objects are subsets of a larger class. For example, a child in the concrete operational stage is able to understand that all roses are flowers, but not all flowers are roses Borst et al. (2013); Politzer (2016). This concept is essential for one’s systematic and logical organizations of conceptual knowledge. Transitivity refers to the ability to understand logical sequences and relationships between objects (Andrews & Halford, 1998; Wright & Smailes, 2015). For instance, if a child knows that Stick A is longer than Stick B, and Stick B is longer than Stick C, they can deduce that Stick A is longer than Stick C.
@@
Mechanical reasoning refers to the ability to understand and apply mechanical concepts and logical principles to solve problems (Allen et al., 2020). This cognitive concept first involves the ability to interpret and predict the behaviors of complex physical systems and understand how different mechanisms of the systems work. Second, mechanical reasoning requires the ability to apply logic rules (O’Brien & Shapiro, 1968; Cesana-Arlotti et al., 2018), such as induction, abduction, syllogism, Modus Ponens and Modus Tollens, and reasoning forms (Byrne, 2016), such as hypotheticals and counterfactuals, figure out how to manipulate these systems to achieve a desired outcome (Hegarty, 2004).
@@
Tool using refers to the ability to manipulate objects in its environment as aids in achieving a specific goal, such as obtaining food or modifying the surroundings. A lot of cognitive components involved in tool using ability, such as affordances, referring to computing the action possibilities offered to the agent by the tool with reference to the agent sensorimotor capabilities (Gibson, 1979). For example, a door handle affords pulling or pushing, indicating how the door should be operated.
@@ 3.4 Data Source
CogDevelop2K comprises 2517 images and 445 videos with multimodal options and questions, crawled primarily from networks as well as self-recorded content. For example, concept intentionality were collected from platforms, including Wikipedia, Reddit, Twitter, Quora, and TieBa. Some of the options were adapted from user comments to ensure content diversity and relevance. Videos for intuitive physics were either self-recorded or produced using Physion [^1] https://physion.net/.
@@ 3.5 Dataset Design
Existing datasets typically support only one question-answer format or single modality type, which hinders the assessment of reasoning capabilities across different modalities within the same domain. For instance, current interleaved image understanding and video understanding models cannot be effectively compared on the same question. To address this limitation, we propose the CogDevelop2k benchmark, which includes multiple Q&A formats (e.g., multiple-choice, true/false, and numeric question-answer) and complicate question-answering by incorporating a new image-video-text interleave format as shown in Fig. 2.
To further explore the cognitive development capabilities of models across these modalities, we optimized CogDevelop2k as follows:
Addressing Weak Image-Text Correlation and Imbalance
In existing interleaved image-text datasets, the correspondence between images and text is often loose, and text provides marginal information for image modeling. This imbalance can cause models to over-rely on textual information, especially when text segments are lengthy (Lin et al. (2023)) . To address this issue and focus on the image understanding abilities of the model, we eliminated sentences that describe the image, such as ’A. an oil paint ¡img¿” This ensures that the textual information is highly relevant to but does not overlap the image content.
Multi-frame questions in CogDevelop can simultaneously evaluate a model’s three inference ability: Co-Reference, Reasoning, and Temporal Understanding (Jiang et al. (2024)). Co-reference involves linking natural language descriptions with specific image inputs (e.g., ”the first image” or ”A.$<$img 1$>$”). Reasoning requires models to make decisions based on cognitive knowledge, such as describing spatial relationships. Temporal Understanding, on the other hand, tests the model’s capability to comprehend sequences of frames in terms of temporal order (multi-frame) and correlation (multi-view) (Li et al. (2024b)). Existing interleaved multi-image datasets can not adequately test all three properties simultaneously. For example, video datasets with temporal information often include only a single video, while multi-image datasets that require co-reference lack temporal dependencies. To address this, CogDevelop introduces multi-video interleaving and video-image interleave formats (multi-frame) to evaluate all three properties concurrently. The statistics of the dataset are presented in Table 2.
@@ 3.6 Evaluation Strategy
We recruited 22 participants, all of whom are Chinese college students proficient in English. Each annotator was asked to label 2 to 6 concepts, with each concept being annotated by two or more annotators. Participants were instructed to skip a question if the question is worded ambiguously or is too complicated to answer in 90 seconds.
@@
Similar to Lu et al. (2022), the zero-shot setup follows the format of $Q(M)T\rightarrow A$, where the input includes the question text (Q), task description (T), and multiple options (M) concatenated as tokens, with the output being the predicted answer (A). Given that model predictions can exhibit bias in multiple-choice settings, we implemented circular evaluation as baseline. In circular evaluation, all answer options are shifted one position at a time, ensuring that the correct answer appears in each option slot. Only when the model correctly predicts all shifted answers is it considered accurate. All images and videos were resized to $488^{2}$. (Liu et al. (2023)).
@@
Strategically crafted prompts can enhance model performance, regardless of whether fine-tuning is applied (Bsharat et al. (2023); Yang et al. (2023)). In contrast to Science Q&A datasets, where image captions are incorporated as the context in the prompts, this approach can cause models to over-rely on text rather than reasoning about the image content. To mitigate this, we use image-independent contexts, such as relevant concept introductions and character assignments, which encourage models to reason beyond the provided textual information. The prompts we used can be categorized into leading words, deeper thing, role assignment, reward/penalty, and explanation. Detailed results can be found in Sec 2.
4 Results and Discussion
@@ 4.1 Cognitive Discussions
Our finding complement earlier research which raises worry that large language models may be ”stochastic parrots” that merely link words and sentences together based on probabilistics but do not understand meanings and logic (Searle, 1980; Bender et al., 2021). If an intelligent agent truly understand that changes in spatial arrangement do not affect quantity, it is logically impossible for it to count correctly the amount of coins when the transformation is shown, while count wrongly when the transformation is not shown (Fig. 6). Contradictions like this reveal that MLLMs virtually do not understand the answers they produce when tackling cognitive reasoning tasks. If this is indeed the case, it may account for a variety of difficulties that MLLMs encounter, particularly in achieving robustness across changing task situations (Zhao et al., 2024).
The developmental trajectory of human cognition is marked by complex cognitive abilities being grounded upon extremely robust understandings of a series of foundational concepts, namely core knowledge (Spelke, 2000; Spelke & Kinzler, 2007). Through early stages of development, children exhibit rudimentary yet stable understandings of objects, actions, number, space, and social partners, each dimension lays the foundations for the acquisitions of complex abilities in later life. It has been suggested that core knowledge is precisely what supports the robustness of human cognition instantiated in commonsense reasoning (Mitchell, 2021). In reverse, the inability to implement core knowledge in artificial intelligence models prevent them from achieving human-level robustness in performances, even if such models seem to excel at certain complex cognitive reasoning tasks (Mitchell, 2020; 2021; Shiffrin & Mitchell, 2023; Palmarini & Mitchell, 2024). MLLMs’ poor performances on foundational concepts like spatiality, permanence, continuity, and perspective, those that directly reflect upon grasps of core knowledge, while achieve proficiency in complex concepts like tool using and intention understanding exactly exemplifies this concern.

### v2

@@ Abstract
Are Multi-modal Large Language Models (MLLMs) stochastic parrots? Do they genuinely understand? This paper aims to explore the core cognitive abilities that human intelligence builds upon to perceive, comprehend, and reason in MLLMs. To this end, we propose CogDevelop2K, a comprehensive benchmark that spans 12 sub-concepts from primitive knowledge like object permanence and boundary to more complex abilities like intentionality understanding, structured via the developmental trajectory of a human mind. We evaluate 46 MLLMs on our benchmarks. Surprisingly, we observe a reversed cognitive developmental trajectory compared to humans. Comprehensively, we further evaluate the influence of evaluation strategies and prompting techniques.
@@ 1 Introduction
Building on the foundation of advanced large language models (LLMs), multi-modal large language models (MLLMs) have recently demonstrated human-level performance in complex tasks involving high-level reasoning, perception, and cognition (Li et al., 2024a; Liu et al., 2024; Gemini, 2023; Fu et al., 2023; OpenAI, 2023), such as Spatial Reasoning (Chen et al., 2024; Cai et al., 2024), OCR (Mori et al., 1999), Scene Understanding (Cordts et al., 2016; Chen et al., 2017), Action Recognition (Jhuang et al., 2013; Herath et al., 2017) and Prediction (Lan et al., 2014; Kong & Fu, 2022). The progress in MLLMs has reignited hopes for achieving Artificial General Intelligence (AGI). However, we pose a critical question: Do MLLMs truly comprehend these tasks and possess the genuine capabilities to perform them, or are they merely ”stochastic parrots” that rely on learning spurious correlations? To explore this, we draw inspiration from the development of human cognition.
@@
As highlighted in previous research, core cognitive abilities form the foundational elements of higher-level human intelligence that existing MLLMs excel at, but such excellency do not translate into a more abstracted, general domain (Mitchell, 2020; 2021; Shiffrin & Mitchell, 2023). The performance of MLLMs on core cognitive tasks therefore provides a more profound insight into their true capacities for knowledge, reasoning, and perception. This, in turn, serves as a critical indicator of whether MLLMs possess genuine intelligence or if they are merely ”stochastic parrots” that depend on learning spurious correlations. To this end, we draw on theoretical and empirical approaches from developmental science to create benchmarks that evaluate core cognitive abilities in large vision-language models, examining the relationships between these abilities. On a high level, we follow Jean Piaget’s theory of cognitive development, which identifies four stages in children: sensorimotor, preoperational, concrete operational, and formal operational (Piaget, 1950; Piaget & Inhelder, 1969; 1974). During the sensorimotor stage, infants acquire knowledge through sensory experiences and actions, developing an understanding of basic object properties, such as permanence, continuity, and boundaries. In the preoperational stage, symbolic representation emerges, along with a grasp of basic physical properties. The concrete operational stage is characterized by the development of logical thinking and an understanding of intuitive physics. Finally, the formal operational stage introduces more advanced cognitive abilities, including abstraction, hypothetical reasoning, counterfactual thinking, and tool use. The interdependence and developmental trajectories of these abilities can be mapped in terms of a tree-like structure (as illustrated in Fig. 1).
To evaluate the performance of MLLMs on the core cognitive abilities, we curate the first-ever vision cognitive development benchmark, termed as CogDeveop2K, which consists of a total of 2519 questions with 2517 images and 455 videos. Then, we evaluate 46 MLLM models on our benchmark that spans all four stages of cognitive development. We introduce a novel multi-frame question format to evaluate models’ co-reference, cognitive reasoning and temporal understanding capability simultaneously. Forty-seven models are compared against a human baseline under zero-shot conditions using 11 different prompts (including no prompt). Surprisingly, while prompts can boost model performance by 8.1%, models still demonstrate reversed trends in cognitive development against those observed in children.
@@ 2.1 Multi-modal Large Language Models
The Vision Language Model (VLM) has a long history from Convolution Neural Networks (CNN) and Recurrent Neural Networks (RNN) (Karpathy & Fei-Fei, 2014; Vinyals et al., 2015) to unified modeling of visual and text modality with transformers (Li et al., 2019; Xu et al., 2023; Tan & Bansal, 2019; Alayrac et al., 2022; Radford et al., 2021). With the advancement of Large Language Models (LLMs), existing state-of-the-art MLLMs (Liu et al., 2024; Li et al., 2023) adopt open-sourced Large Language Models such as Llama (Touvron et al., 2023), Mistral (Jiang et al., 2023). Instruction tuning is also introduced to further improve the task generalization ability of MLLMs (Liu et al., 2024; Dai et al., 2023). To acquire open-ended conversation abilities, LLaVA (Liu et al., 2024) proposes to distill the conversational abilities of ChatGPT to MLLMs, boosting performance by a large margin, which becomes a defacto procedure in the area (Wang et al., 2023; Bai et al., 2023; Gemini, 2023; Team, 2024; Sun et al., 2023; Li et al., 2022).
@@ 2.2 Human Cognitive Development
Preoperational and concrete operational stage are the second and third stage of Piaget’s cognitive development. Typically spanning over 2 to 7 years of age, preoperational stage is the transitional stage to concrete operational stage, which children enter around 7 years of age. During this period, children begin to develop internalized mental actions supported by organized structures that can be manipulated and reversed in systematic ways, known as mental operations (Janet, 1905; Kirkpatrick, 1908; Piaget, 1950; Piaget & Inhelder, 2014; Miller, 2016). Through mental operations, children are then able to rigidly perform tasks that are previously unreachable, such as thinking from other people’s perspectives, understanding hierarchical relations of objects, and reasoning about physical events in the world. These tasks require not only rudimentary understandings of physical concepts, which gradually became in place during preoperational stage, but also relational and transformational reasoning that can only be done through mental operations (Piaget & Inhelder, 1974; Church & Goldin-Meadow, 1986; Houdé, 1997). Since preoperational stage is mostly meaningful as the transitional period preceding concrete operational stage, we do not have evaluation dimensions specifically targeting the stage. However, tasks targeting concrete operational stage could assess the existence of knowledge associated with preoperational stage, such as law of conservation (Piaget, 1952; Halford, 2011; Houdé, 1997).
@@ 2.3 Evaluation Dimension
Boundary refers to the cognitive understanding of where one object ends and another begins, an essential aspect of perceiving and understanding the physical world (Kestenbaum et al., 1987). Without understanding boundary, it seems very hard to construct a concept of object (Berkeley, 1709; Jackendoff, 1991).
@@
Permanence, or specifically object permanence, is the idea that objects continue to exist even when they are not visible (Baillargeon, 1986; Spelke et al., 1992). Imagine a simple scene: a small child playing peek-a-boo. In the beginning, when the caregiver covers their face with their hands, the child might seem surprised or even distressed, thinking the person has disappeared. However, as children’s understanding of permanence develops, they begin to realize that just because they can’t see the person’s face, it doesn’t mean the person is gone.
@@
Hierarchical relation refers to the ability to organize objects or concepts into structured categories and subcategories, which are supported by the development of mental operations marked by class inclusion and transitivity (Shipley, 1979; Winer, 1980; Chapman & McBride, 1992). Class inclusion refers to the ability to recognize that some classes or groups of objects are subsets of a larger class. For example, a child in the concrete operational stage is able to understand that all roses are flowers, but not all flowers are roses (Borst et al., 2013; Politzer, 2016). This concept is essential for one’s systematic and logical organizations of conceptual knowledge. Transitivity refers to the ability to understand logical sequences and relationships between objects (Andrews & Halford, 1998; Wright & Smailes, 2015). For instance, if a child knows that Stick A is longer than Stick B, and Stick B is longer than Stick C, they can deduce that Stick A is longer than Stick C.
@@
Mechanical reasoning refers to the ability to understand and apply mechanical concepts and logical principles to solve problems (Allen et al., 2020). This cognitive concept first involves the ability to interpret and predict the behaviors of complex physical systems and understand how different mechanisms of the systems work. Second, mechanical reasoning requires the ability to apply logic rules, such as induction, abduction, syllogism (O’Brien & Shapiro, 1968; Cesana-Arlotti et al., 2018), and reasoning forms, such as hypotheticals and counterfactual (Byrne, 2016), to figure out how to manipulate these systems to achieve a desired outcome (Hegarty, 2004).
@@
Tool using refers to the ability to manipulate tool objects in its environment as aids in achieving a specific goal, such as obtaining food or modifying the surroundings. A lot of cognitive components involved in tool using ability, such as affordances, referring to computing the action possibilities offered to the agent by the tool with reference to the agent’s sensorimotor capabilities (Gibson, 1979). For example, a door handle affords pulling or pushing, as how the door should be operated by a human agent.
@@ 2.4 Data Source
CogDevelop2K comprises 2517 images and 445 videos with multimodal options and questions, crawled primarily from networks as well as self-recorded content. Our targeted platforms include Wikipedia, Reddit, Twitter, Quora, and TieBa. Some of the captions were adapted from user comments to ensure content diversity and relevance. Videos for intuitive physics were either self-recorded or produced using Physion [^1] https://physion.net/.
@@ 2.5 Dataset Design
Existing datasets typically support only one question-answer format or single modality type, which hinders the assessment of reasoning capabilities across different modalities within the same domain. For instance, current interleaved image understanding and video understanding models cannot be effectively compared on the same question. To address this limitation, we include multiple Q&A formats (e.g., multiple-choice, true/false, and numeric question-answer) and complicate question-answering by incorporating a new image-video-text interleave format as shown in Fig. 2. To further explore the cognitive development capabilities of models across these modalities, we optimize our formality as follows:
Addressing Weak Image-Text Correlation and Imbalance
In existing interleaved image-text datasets, the correspondence between images and text is often loose, and text provides marginal information for image modeling. This imbalance can cause models to over-rely on textual information, especially when text segments are lengthy (Lin et al., 2023) . To address this issue and focus on the image understanding abilities of the model, we eliminate sentences that describe the image. This ensures that the textual information is highly relevant to but does not overlap with the image content.
Multi-frame questions can simultaneously evaluate a model’s three inference ability: Co-Reference, Reasoning, and Temporal Understanding (Jiang et al., 2024). Co-reference involves linking natural language descriptions with specific image inputs (e.g., ”the first image” or ”A.$<$img 1$>$”). Reasoning requires models to make decisions based on cognitive knowledge, such as describing spatial relationships. Temporal Understanding, on the other hand, tests the model’s capability to comprehend sequences of frames in terms of temporal order (multi-frame) and correlation (multi-view) (Li et al., 2024b). Existing interleaved multi-image datasets can not adequately test all three properties simultaneously. For example, video datasets with temporal information often include only a single video, while multi-image datasets that require co-reference lack temporal dependencies. To address this, we introduce multi-video interleaving and video-image interleave formats (multi-frame) to evaluate all three properties concurrently. The statistics of the dataset are presented in Table 1.
@@ 2.6 Evaluation Strategy
We recruit 22 participants, all of whom are college students proficient in English. Participants are instructed to skip a question if the question is phrased ambiguously or is too complicated to answer in 90 seconds. This question is marked as failed if the human participant does not provide an answer.
@@
Similar to previous studies (Lu et al., 2022), the zero-shot setup follows the format of $Q(M)T\rightarrow A$, where the input includes the question text (Q), task description (T), and multiple options (M) concatenated as tokens, with the output being the predicted answer (A). Given that model predictions can exhibit bias in multiple-choice settings, we implemented circular evaluation as baseline. In circular evaluation, all answer options are shifted one position at a time, ensuring that the correct answer appears in each option slot. Only when the model correctly predicts all shifted answers is it considered accurate (Liu et al., 2023). All images and videos were resized to $488^{2}$.
@@
Strategically crafted prompts can enhance model performance, regardless of whether fine-tuning is applied (Bsharat et al., 2023; Yang et al., 2023). To mitigate this, we use image-independent contexts, such as relevant concept introductions and character assignments, which encourage models to reason beyond the provided textual information. The prompts we used can be categorized into leading words, deeper thing, role assignment, reward or penalty, and explanation.
3 Results
@@ 4 Discussion
Our findings complement earlier research which raises worry that large language models may be ”stochastic parrots” that merely link words and sentences together based on probabilities but do not understand meanings and logic (Searle, 1980; Bender et al., 2021). If an intelligent agent truly understand that changes in spatial arrangement do not affect quantity, it is logically impossible for it to count correctly the amount of coins when the transformation is shown, while count wrongly when the transformation is not shown (Fig. 6). Contradictions like this reveal that MLLMs virtually do not understand the answers they produce when tackling cognitive reasoning tasks. If this is indeed the case, it may account for a variety of difficulties that MLLMs encounter, particularly in achieving robustness across changing task situations (Zhao et al., 2024).
The developmental trajectory of human cognition is marked by complex cognitive abilities being grounded upon extremely robust understandings of a series of foundational concepts, namely core knowledge (Spelke, 2000; Spelke & Kinzler, 2007). Through early stages of development, children exhibit rudimentary yet stable understandings of objects, actions, number, space, and social partners, each dimension laying the foundations for the acquisitions of complex abilities in later life. It has been suggested that core knowledge is precisely what supports the robustness of human cognition instantiated in commonsense reasoning (Mitchell, 2021). In reverse, the inability to implement core knowledge in artificial intelligence models prevent them from achieving human-level robustness in performances, even if such models seem to excel at certain complex cognitive reasoning tasks (Mitchell, 2020; 2021; Shiffrin & Mitchell, 2023; Palmarini & Mitchell, 2024). MLLMs’ poor performances on foundational concepts like spatiality, permanence, continuity, and perspective-taking, which directly reflect upon grasps of core knowledge, while achieve proficiency in complex concepts like tool using and intention understanding exactly exemplifies this concern.